# Improving Lesion Segmentation in Medical Images by Global and Regional Feature Compensation

### v1

@@ Abstract
Automated lesion segmentation of medical images has made tremendous improvements in recent years due to deep learning advancements. However, accurately capturing fine-grained global and regional feature representations remains a challenge. Many existing methods obtain suboptimal performance on complex lesion segmentation due to information loss during typical downsampling operations and the insufficient capture of either regional or global features. To address these issues, we propose the Global and Regional Compensation Segmentation Framework (GRCSF), which introduces two key innovations: the Global Compensation Unit (GCU) and the Region Compensation Unit (RCU). The proposed GCU addresses resolution loss in the U-shaped backbone by preserving global contextual features and fine-grained details during multiscale downsampling. Meanwhile, the RCU introduces a self-supervised learning (SSL) residual map generated by Masked Autoencoders (MAE), obtained as pixel-wise differences between reconstructed and original images, to highlight regions with potential lesions. These SSL residual maps guide precise lesion localization and segmentation through a patch-based cross-attention mechanism that integrates regional spatial and pixel-level features. Additionally, the RCU incorporates patch-level importance scoring to enhance feature fusion by leveraging global spatial information from the backbone. Experiments on two publicly available medical image segmentation datasets, including brain stroke lesion and coronary artery calcification datasets, demonstrate that our GRCSF outperforms state-of-the-art methods, confirming its effectiveness across diverse lesion types and its potential as a generalizable lesion segmentation solution.
@@
Medical Image Processing , Lesion Segmentation , Global and Regional Feature , Representation Learning , Self-supervised Learning , Transformer , Mask Autoencoders
@@ 1 Introduction
In medical imaging, accurate lesion segmentation plays a critical role in assisting clinicians with diagnosing diseases, monitoring disease progression, and evaluating treatment effectiveness [1]. Despite advancements in automated segmentation methods, accurately localizing and delineating lesions with complex characteristics remains challenging [2]. For example, ischemic stroke lesions in the brain pose a significant segmentation challenge due to their low-contrast, where the pixel intensities are similar to those of the surrounding tissues, referred to as isointense [3] in MR imaging. The limitations of T1-weighted imaging to capture edema and early ischemic changes [4] make low-contrast stroke lesions difficult to distinguish from normal tissues. These lesions often exhibit irregular and ambiguous boundaries [5], further complicating their segmentation using traditional deep learning methods. Another example of challenging segmentation task is coronary artery calcification (CACS), occupying only a small region of the image, often measuring less than 5 mm in length and corresponding to 10–13 pixels referring to different in-plane resolution in the non-contrast CT slices used in this study [6]. The small size and irregular shape of the CACS, along with its variability in location and density, make precise segmentation and quantification difficult. Furthermore, non-contrast CT, the primary imaging modality for detecting and assessing CACS [7], further complicates the process due to its low signal-to-noise ratio, resulting in blurred boundaries and imprecise delineation.
@@
Automated segmentation methods have improved significantly in recent years, driven by advances in deep learning [8, 9]. Convolutional neural networks (CNN) [10], particularly encoder-decoder architectures [11] such as U-Net [12] and UNet ++ [13], have been widely applied for medical image segmentation task. These architectures preserve regional context but failed to capture global context due to fixed-size convolutional kernels, limiting their ability to segment lesions with varying locations or ambiguous boundaries. To address these limitations, advanced CNN-based models, such as DeepLabv3 [14] implement various receptive fields to improve the segmentation of complex anatomical structures, but remain limited by the fundamental constraints of convolutional operations, particularly in modeling long-range dependencies.
@@
In recent years, transformer-based models [16, 17, 18, 19, 20] have demonstrated their potential in addressing these limitations by modeling global context and capturing long-range dependencies through self-attention mechanisms. TransUNet [17] and SwinUNet [19] integrate hybrid architectures to learn regional and global characteristics, resulting in improved segmentation performance. However, they rely on skip connections to transfer low-level features from the encoder to the decoder for spatial detail recovery. This approach may not fully capture fine-grained details, limiting their localization ability and accuracy in segmenting challenging lesions. Several other transformer-based models [21, 22, 23] have been developed to solve the challenges of medical image segmentation. While these models have shown notable improvements in handling complex and variable structures, they still face challenges in accurately localizing subtle and irregular lesions and delineating their boundaries. This limitation arises because patch-based self-attention tends to blur regional details and introduce feature inconsistencies among patches. SPiN [24] is developed to address these challenges using the subpixel mechanism. Although effective in handling small stroke lesion segmentation, SPiN relies heavily on supervised learning approach and requires extensive manually annotated datasets.
@@
Recent methods have enhanced performance in segmentation tasks by employing self-supervised learning (SSL) [25], which learns meaningful representations from unlabeled data. SSL reduces reliance on costly and time-intensive manual annotations while improving downstream task performance. It achieves this through pretext tasks that exploit the inherent structure of images, such as contrastive learning [26, 27, 28], pixel-level image reconstruction [29], and masked token prediction [30, 31]. These tasks enable models to extract robust global features [32, 33] from datasets and localized features from individual images, which can serve as initialization weights for supervised learning. SSL has demonstrated particular effectiveness in medical imaging [34, 35, 36], showing that SSL pre-training improves segmentation performance, especially when fine-tuned with small annotated datasets. However, many SSL methods focus predominantly on learning high-level global features that capture the overall data context, often overlooking fine-grained, pixel-level details. These details, critical for segmentation tasks, such as precisely delineating the boundaries of challenging lesions with low-contrast or small sizes, remain unexplored in existing SSL approaches. This gap underscores the need for SSL methods that integrate global and pixel-level information to enhance segmentation performance. Masked Autoencoders (MAE) [29], introduced by He et al., addresses this limitation by incorporating localized feature representations. Although MAE has been used primarily as a pre-training technique in medical imaging [37, 38], its potential for tasks beyond pre-training, such as using reconstructed images to guide segmentation remains unexplored.
@@
To address the limitations mentioned above, we propose a novel dual-feature compensation framework, named Global and Regional Compensation Segmentation Framework (GRCSF), to improve the medical image segmentation of challenging lesions. GRCSF adopts an encoder-decoder structure built on UNet++ and takes the raw images as input. It compensates for the detailed global features that are often missing in CNN-based models. Unlike traditional approaches that use MAE for pre-training, our method, to our knowledge, is the first medical image segmentation framework to use MAE’s reconstruction process to generate SSL residual maps that capture pixel-level differences between the original images and their reconstructions. Therefore, it mitigates the inconsistency between global and local feature representations introduced by transformer models. The key contributions of our work are as follows:
@@
    We introduce the Regional Compensation Unit (RCU) module within the decoder, by applying a cross-attention mechanism on residual maps learned from the SSL strategy and the features of UNet++. We can obtain importance score maps to better highlight those regions with a higher likelihood of containing lesions by incorporating additional global features from the backbone model, which helps reduce false-positive regions on the SSL residual maps. The RCU addresses the challenge of localizing blurred lesions that are often difficult to detect based on limited regional spatial features.
@@
    We validate the effectiveness and robustness of GRCSF for general lesion segmentation, using two representative datasets: ATLAS 2.0, which consists of low-contrast brain stroke lesions, and orCaScore, which contains small-sized coronary artery calcifications.
@@ 2.1 Deep Learning for Supervised Medical Image Segmentation
The CNN-based encoder-decoder architectures [11] such as U-Net [12] and its variant UNet ++ [13], have been widely adopted for the medical image segmentation task. These architectures utilize skip connections to preserve spatial resolution, making them effective for various segmentation tasks. Unet++ improved multi-scale feature extraction by introducing nested skip pathways, further enhancing its ability to handle complex and heterogeneous medical data. DeepLabv3 [14] introduced Atrous Spatial Pyramid Pooling (ASPP) [15] to encode multi-scale contextual information by probing features at multiple rates and fields of view. A recent study SPiN [24] has made efforts to accurately segment isointense brain strok lesions. It developed a subpixel embedding mechanism to generate high-resolution confidence maps and employed a learnable downsampler to accurately localize irregular lesions and delineate boundaries. While CNN-based segmentation models have demonstrated improvement in medical image segmentation, they are still limited to capture global features due to fixed receptive field.
Transformer-based models TransUNet [17] combined CNNs and transformers in a hybrid architecture, leveraging CNNs for high-resolution spatial encoding and transformers for capturing long-range dependencies. Similarly, SwinUNet [19] adopted a fully transformer-based U-shaped model, incorporating a hierarchical Swin Transformer [20] with a shifted windows mechanism as the encoder to simultaneously model global and local dependencies. However, the use of window-based attention in transformer-based models can introduce boundary discontinuities, which is particularly problematic for lesions with ambiguous edges. Moreover, its performance is heavily dependent on large training datasets and extensive parameter tuning, making adaptation to new datasets challenging.
Existing methods either fail to effectively capture global features or lack consistency in regional feature representation. GRCSF utilizes pixel-level similarity from different feature scales to enhance feature representations.
@@ 2.2 Self-Supervised Learning for Medical Image Segmentation
SSL leverages unlabeled data to learn meaningful feature representations for different tasks. For example, MAE, originally developed for natural images, have also shown potential in medical imaging by reconstructing masked image patches and learning global semantic and localized contextual features. MAE employs an encoder-decoder architecture where a portion of image patches is randomly masked during training. The encoder extracts semantic information from visible patches, while the decoder reconstructs the masked regions. This approach allows MAE to learn the overall structure of the dataset and the unique fine-grained details of each image. Zhou et al. [37] has shown that MAE self pre-training improves various medical image tasks, including abdominal CT multi-organ segmentation, and MRI brain tumor segmentation. Mask in Mask (MiM) [38] framework introduced hierarchical token learning and cross-level alignment mechanisms to improve downstream segmentation tasks across large-scale datasets.
These advances highlight the growing impact of SSL in medical image segmentation, enabling models to leverage unlabeled data effectively and generalize across diverse medical imaging tasks. However, most SSL methods remain primarily focused on pre-training. To our best knowledge, GRCSF is the first medical image segmentation work to exploit the capacity of SSL to recover global features and produce residual maps that highlight the regions of potential lesion used for regional feature compensation. This operation takes advantage of additional features to improve the likelihood at the pixel-level of accurately identifying abnormal regions.
@@ 3.2 Global Compensation Unit
[…]
The skip-concatenation residual map is defined as:
[…]
@@
[…]
Here, $\text{SE}(\mathit{U})$ and $\text{SE}(\mathit{F})$ are attention maps generated by the SE block. $\otimes$ and $\oplus$ denote the element-wise product and addition, respectively. Each position can capture contextual information in the horizontal and vertical directions. The residual map $R$ highlights areas of significant change and is particularly relevant for architectures such as UNet++, where each layer involves multiple skip concatenation features. For simpler architectures such as UNet, the residual map $R$ is not required, as the updated skip feature $C$ alone is sufficient to perform skip concatenation between the encoder and decoder.
@@ 3.3.1 Residual Maps from Self-Supervised Learning
The SSL residual maps $RM_{1}$ and $RM_{2}$ are computed by calculating the cosine similarity pixel-wise between the input image $\mathit{I}$ and the averaged reconstructions $R_{1}$ and $R_{2}$, respectively, and subtracting the result from 1. The pixel-wise cosine similarity between two images is defined as CS:
$\text{CS}(I,R)=\frac{\sum_{x=1}^{H}\sum_{y=1}^{W}I(x,y)R(x,y)}{\sqrt{\sum_{x=1}^{H}\sum_{y=1}^{W}I(x,y)^{2}}\cdot\sqrt{\sum_{x=1}^{H}\sum_{y=1}^{W}R(x,y)^{2}}}$
[…]
where (x, y) is pixel coordinates. This metric measures the cosine of the angle between the two vectors, quantifying their similarity on a scale from -1 to 1, where 1 indicates perfect alignment and -1 indicates complete opposition. These MRM highlight regions where the reconstructions differ from the original image, reflecting potential lesion areas (Fig. 3).

### v2

@@ Abstract
Automated lesion segmentation of medical images has made tremendous improvements in recent years due to deep learning advancements. However, accurately capturing fine-grained global and regional feature representations remains a challenge. Many existing methods achieve suboptimal performance in complex lesion segmentation due to information loss during typical downsampling operations and insufficient capture of either regional or global features. To address these issues, we propose the Global and Regional Compensation Segmentation Framework (GRCSF), which introduces two key innovations: the Global Compensation Unit (GCU) and the Region Compensation Unit (RCU). The proposed GCU addresses resolution loss in the U-shaped backbone by preserving global contextual features and fine-grained details during multiscale downsampling. Meanwhile, the RCU introduces a self-supervised learning (SSL) residual map generated by Masked Autoencoders (MAE), obtained as pixel-wise differences between reconstructed and original images, to highlight regions with potential lesions. These SSL residual maps guide precise lesion localization and segmentation through a patch-based cross-attention mechanism that integrates regional spatial and pixel-level features. Additionally, the RCU incorporates patch-level importance scoring to enhance feature fusion by leveraging global spatial information from the backbone. Experiments on three publicly available medical image segmentation datasets, including brain stroke lesion, lung tumor and coronary artery calcification datasets, demonstrate that our GRCSF outperforms state-of-the-art methods, confirming its effectiveness across diverse lesion types and its potential as a generalizable lesion segmentation solution.
@@
Medical Image Processing , Lesion Segmentation , Global and Regional Feature , Representation Learning , Self-supervised Learning , Mask Autoencoders
@@ 1 Introduction
In medical imaging, accurate lesion segmentation plays a critical role in assisting clinicians with diagnosing diseases, monitoring disease progression, and evaluating treatment effectiveness [1]. Despite advancements in automated segmentation methods, accurately localizing and delineating lesions with complex characteristics remains challenging. For example, ischemic stroke lesions in the brain pose a significant segmentation challenge due to their low-contrast, where the pixel intensities are similar to those of the surrounding tissues, referred to as isointense [2] in MR imaging. The limitations of T1-weighted imaging to capture edema and early ischemic changes [3] make low-contrast stroke lesions difficult to distinguish from normal tissues. Similar to brain lesions, non-small-cell lung cancer (NSCLC) tumors on thin-section chest CT, despite their generally higher contrast, remain difficult to delineate accurately because their shapes, sizes, and locations vary markedly across patients. These lesions often exhibit irregular and ambiguous boundaries, further complicating their segmentation with traditional deep learning methods. Another example of challenging segmentation task is coronary artery calcification (CACS), occupying only a small region of the image, often measuring less than 5 mm in length and corresponding to 10–13 pixels referring to different in-plane resolution in the non-contrast CT slices used in this study [4]. The small size and irregular shape of the CACS, along with its variability in location and density, make precise segmentation and quantification difficult. Furthermore, non-contrast CT, the primary imaging modality for detecting and assessing CACS, further complicates the process due to its low signal-to-noise ratio, resulting in blurred boundaries and imprecise delineation.
@@
Automated segmentation methods have improved significantly in recent years, driven by advances in deep learning [5]. Convolutional neural networks (CNN), particularly encoder-decoder architectures such as U-Net [6] and UNet ++ [7], have been widely applied for medical image segmentation task. These architectures preserve regional context but failed to capture global context due to fixed-size convolutional kernels, limiting their ability to segment lesions with varying locations or ambiguous boundaries. To address these limitations, advanced CNN-based models [8, 9, 10], such as DeepLabv3 [11] implement various receptive fields to improve the segmentation of complex anatomical structures, but remain limited by the fundamental constraints of convolutional operations, particularly in modeling long-range dependencies.
@@
In recent years, transformer-based models have demonstrated their potential in addressing these limitations by modeling global context and capturing long-range dependencies through self-attention mechanisms. TransUNet [12], TransFuse [13], UTNet [14] and SwinUNet [15] integrate hybrid architectures to learn regional and global characteristics, resulting in improved segmentation performance. However, they rely on skip connections to transfer low-level features from the encoder to the decoder for spatial detail recovery. This approach may not fully capture fine-grained details, limiting their localization ability and accuracy in segmenting challenging lesions. Several other transformer-based models [16] have been developed to solve the challenges of medical image segmentation. While these models have shown notable improvements in handling complex and variable structures, they still face challenges in accurately localizing subtle and irregular lesions and delineating their boundaries. This limitation arises because patch-based self-attention tends to blur regional details and introduce feature inconsistencies among patches. SPiN [17] is developed to address these challenges using the subpixel mechanism. Although effective in handling small stroke lesion segmentation, SPiN relies heavily on supervised learning approach and requires extensive manually annotated datasets.
@@
Recent methods have enhanced performance in segmentation tasks by employing self-supervised learning (SSL), which learns meaningful representations from unlabeled data. SSL reduces reliance on costly and time-intensive manual annotations while improving downstream task performance. It achieves this through pretext tasks that exploit the inherent structure of images, such as contrastive learning [18], pixel-level image reconstruction [19, 20], and masked token prediction [21]. These tasks enable models to extract robust global features [22] from datasets and localized features from individual images, which can serve as initialization weights for supervised learning. SSL has demonstrated particular effectiveness in medical imaging [23], showing that SSL pre-training improves segmentation performance, especially when fine-tuned with small annotated datasets. However, many SSL methods focus predominantly on learning high-level global features that capture the overall data context, often overlooking fine-grained, pixel-level details. These details, critical for segmentation tasks, such as precisely delineating the boundaries of challenging lesions with low-contrast or small sizes, remain unexplored in existing SSL approaches. This gap underscores the need for SSL methods that integrate global and pixel-level information to enhance segmentation performance. Masked Autoencoders (MAE) [19], introduced by He et al., addresses this limitation by incorporating localized feature representations. Although MAE has been used primarily as a pre-training technique in medical imaging [24, 25], its potential for tasks beyond pre-training, such as using reconstructed images to guide segmentation remains unexplored.
Foundation models such as the Segment Anything Model (SAM) [26] and GPT-4 with vision (GPT-4V) [27] have emerged as powerful tools for enabling zero-shot and few-shot segmentation through user-provided prompts. SAM performs segmentation based on prompts such as points, bounding boxes, or masks. Trained on over 1 billion masks across 11 million images, SAM can deliver competitive or even superior performance compared to traditional supervised models in a zero-shot or few-shot scenarios. However, its performance is highly dependent on the quality of the input prompts. Since SAM lacks the ability to correct prompts that entirely fall within false-positive regions, inaccurate or suboptimal prompts can lead to poor segmentation results. GPT-4V, a vision-enhanced large language model (LLM), incorporates image inputs and opens new perspectives for generating image-text pairs. Despite its potential, as it is not specifically trained for the medical domain or for segmentation tasks, GPT-4V may understand general content in medical images but lacks the fine-grained or pixel-level understanding needed to localize complex anatomical structures and lesions [28]. As a result, it cannot reliably generate accurate bounding boxes or masks as prompts for downstream segmentation models, nor can it serve directly as a segmentation model due to the absence of a dedicated segmentation head.
@@
To address the limitations mentioned above, we propose a novel dual-feature compensation framework, named Global and Regional Compensation Segmentation Framework (GRCSF), to improve the medical image segmentation of challenging lesions. GRCSF adopts an encoder-decoder structure built on UNet++ and takes the raw images as input. It compensates for the detailed global features that are often missing in CNN-based models. Unlike traditional approaches that use MAE for pre-training, our method, to our knowledge, is the first medical image segmentation framework to use MAE’s reconstruction process to generate SSL residual maps that capture pixel-level differences between the original images and their reconstructions. Therefore, this mitigates inconsistencies between global and local feature representations introduced by transformer models. The key contributions of our work are as follows:
@@
    We introduce the Regional Compensation Unit (RCU) module within the decoder, by applying a cross-attention mechanism between residual maps learned from the SSL strategy and the features of UNet++. We can obtain importance score maps to better highlight those regions with a higher likelihood of containing lesions by incorporating additional global features from the backbone model, which helps reduce false positives in SSL residual maps. The RCU addresses the challenge of localizing blurred lesions, which are often difficult to detect due to limited regional spatial features.
@@
    We validate the effectiveness and robustness of GRCSF for general lesion segmentation, using three representative datasets: ATLAS 2.0, including low-contrast brain stroke lesions; MSD Lung Tumor Segmentation, featuring irregular NSCLC tumors; and orCaScore, which contains small-sized coronary artery calcifications.
@@ 2.1 Deep Learning for Supervised Medical Image Segmentation
The CNN-based encoder-decoder architectures such as U-Net [6] and its variant UNet ++ [7], have been widely adopted for the medical image segmentation task. These architectures utilize skip connections to preserve spatial resolution, making them effective for various segmentation tasks. Unet++ improved multi-scale feature extraction by introducing nested skip pathways, further enhancing its ability to handle complex and heterogeneous medical data. DeepLabv3 [11] introduced Atrous Spatial Pyramid Pooling (ASPP) [29] to encode multi-scale contextual information by probing features at multiple rates and fields of view. A recent study SPiN [17] has made efforts to accurately segment isointense brain strok lesions. It developed a subpixel embedding mechanism to generate high-resolution confidence maps and employed a learnable downsampler to accurately localize irregular lesions and delineate boundaries. While CNN-based segmentation models have demonstrated improvement in medical image segmentation, they are still limited to capture global features due to fixed receptive field.
Transformer-based models TransUNet [12], TransFuse [13] and UTNet [14] combined CNNs and transformers in a hybrid architecture, leveraging CNNs for high-resolution spatial encoding and transformers for capturing long-range dependencies. Similarly, SwinUNet [15] adopted a fully transformer-based U-shaped model, incorporating a hierarchical Swin Transformer [30] with a shifted windows mechanism as the encoder to simultaneously model global and local dependencies. However, the use of window-based attention in transformer-based models can introduce boundary discontinuities, which is particularly problematic for lesions with ambiguous edges. Moreover, its performance is heavily dependent on large training datasets and extensive parameter tuning, making adaptation to new datasets challenging.
Existing methods either fail to effectively capture global features or lack of consistent regional feature representation. GRCSF utilizes pixel-level similarity from different feature scales to enhance feature representations.
@@ 2.2 Self-Supervised Learning for Medical Image Segmentation
SSL leverages unlabeled data to learn meaningful feature representations for different tasks. For example, MAE, originally developed for natural images, have also shown potential in medical imaging by reconstructing masked image patches and learning global semantic and localized contextual features. MAE employs an encoder-decoder architecture in which a portion of image patches are randomly masked during training. The encoder extracts semantic information from visible patches, while the decoder reconstructs the masked regions. This approach allows MAE to learn the overall structure of the dataset and the unique fine-grained details of each image. Zhou et al. [24] has shown that MAE self pre-training improves various medical image tasks, including abdominal CT multi-organ segmentation, and MRI brain tumor segmentation. Mask in Mask (MiM) [25] framework introduced hierarchical token learning and cross-level alignment mechanisms to improve downstream segmentation tasks across large-scale datasets.
These advances highlight the impact of SSL in medical image segmentation, enabling models to leverage unlabeled data effectively and generalize across diverse medical imaging tasks. However, most SSL methods remain primarily focused on pre-training. To our best knowledge, GRCSF is the first medical image segmentation work to exploit the capacity of SSL to recover global features and produce residual maps that highlight the regions of potential lesion used for regional feature compensation. This operation takes advantage of additional features to improve the likelihood at the pixel-level of accurately identifying abnormal regions.
2.3 Transformer-based Vision Foundation Models for Medical Image Segmentation
Recent developments in large-scale vision foundation models have significantly influenced medical image segmentation. SAM  [26] is a representative prompt-driven segmentation framework and has demonstrated robust zero-shot performance across diverse medical imaging modalities  [31]. However, its effectiveness is highly dependent on the quality of the input prompts. To address this, previous studies  [32] have explored automated prompt generation strategies from coarse segmentation masks to enhance segmentation accuracy.
GPT-4V  [27] has shown its ability in modality recognition, disease diagnosis, medical visual question answering (VQA) and report generation. However, its performance is limited in certain tasks such as chest radiograph interpretation  [28] . Moreover, GPT-4V failed to localize the region-of-interest when it shared similar texture and shape with the background, highlighting its vulnerability to complex backgrounds  [33]. Although integrating such generalist models into clinical segmentation workflows remains a challenge, advances in prompt engineering  [34, 35] offer a potential path toward improving localization accuracy and segmentation robustness.
@@ 3.2 Global Compensation Unit
The GCU addresses pixel-level information loss caused by downsampling in U-shaped CNN architectures by reintroducing lost details into the skip-concatenated features. During downsampling, the input feature map undergoes two convolutional operations, resulting in a downsampled feature map $\mathit{S}\in\mathbb{R}^{H\times W\times C}$. To mitigate information loss, $\mathit{S}$ is re-upsampled to match the resolution of the skip feature map from the previous layer, $\mathit{F}\in\mathbb{R}^{H\times W\times C}$, producing $\mathit{RU}\in\mathbb{R}^{H\times W\times C}$. A comparison between $\mathit{RU}$ and $\mathit{F}$ identifies regions of information loss.
To refine $\mathit{RU}$, it is multiplied by the corresponding upsampled feature map $\mathit{U}\in\mathbb{R}^{H\times W\times C}$ in the decoder, which is enhanced by a Squeeze-and-Excitation (SE) mechanism  [36]. The SE block, applied to both $\mathit{F}$ and $\mathit{U}$, focuses on critical regions while suppressing irrelevant areas. This process yields two key outputs: The skip-concatenation residual map is defined as:
$R=\text{CS}(\mathit{RU}\otimes\text{SE}(U),\text{SE}(F))$
@@
$C=\text{CS}(\mathit{RU}\otimes\text{SE}(U),\text{SE}(F))\otimes\mathit{F}\oplus\mathit{F}$
Here, $\text{SE}(\mathit{U})$ and $\text{SE}(\mathit{F})$ are attention maps generated by the SE block. CS is the pixel-wise cosine similarity operation, while $\otimes$ and $\oplus$ denote the element-wise product and addition, respectively. Each position captures context along both horizontal and vertical axes. The residual map $R$ highlights areas of significant change and is particularly relevant for architectures such as UNet++, where each layer involves multiple skip concatenation features. For simpler architectures such as UNet, the residual map $R$ is not required, as the updated skip feature $C$ alone is sufficient to perform skip concatenation between the encoder and decoder.
@@ 3.3.1 Residual Maps from Self-Supervised Learning
The SSL residual maps $RM_{1}$ and $RM_{2}$ are computed by calculating the pixel-wise absolute difference (AbsDiff) between the input image $\mathit{I}$ and the averaged reconstructions $R_{1}$ and $R_{2}$, respectively. The pixel-wise AbsDiff between two images is defined as:
$\text{AbsDiff}(I,R)=\left|\mathit{I}(x,y)-\mathit{R}(x,y)\right|$
$\mathit{RM}_{1}=\text{AbsDiff}(I,R_{1})$
$\mathit{RM}_{2}=\text{AbsDiff}(I,R_{2})$
where (x, y) are the pixel coordinates. This metric quantifies the AbsDiff between two arrays at corresponding pixel locations, representing their similarity on a scale from 0 to 1, where 0 indicates perfect alignment and 1 indicates complete opposition. These MRM highlight regions where the reconstructions differ from the original image, reflecting potential lesion areas (Fig. 3).